Laser Intervention History
 Share
The version of the browser you are using is no longer supported. Please upgrade to a supported browser.Dismiss

 
£
%
123
 
 
 
 
 
 
 
 
 
 
 
 
 
 
ABCDEFGHIJKLMNOPQRSTUVW
1
# Log file for interventions on the Photonics Laser system of the CMS ECAL
2
3
# Please keep the intervention type in a numeric format and start non-data lines with `#'. It is needed for parsing the file in txt format when producing plots.
4
5
# Flag legend:1=cooling; 2=shielding; 3=B field; 4=software; 5=measurement; 6=hardware repair/cleaned/changed; 7=changed current; 8=changed timing
6
7
# Begin  Date/TimeEnd Date/TimeFlagLaser ON or OFFComments/observations
8
26/11/201226/11/20126ONpower measurement; clean main fibre on laser side B=0T
9
17/09/201221/09/20126ON-OFFTechnical stop. Liyuan. B=0T
10
13/09/201213/09/20126OFFPower measurement. New chiller to add the heater option module next Technical Stop.
11
03/09/201203/09/20125OFFPower measurement
12
28/08/201211:00:0028/08/201215:30:001OFF11 am to ~15 30 pm, Cooler Off: fuse replaced, Laser was off since ~10 30 am. Took time to get good fuse. Laser turn back on without problem
13
28/08/201211:00:0028/08/201215:30:001OFF11 am to ~15 30 pm, Cooler Off: fuse replaced, Laser was off since ~10 30 am. Took time to get good fuse. Laser turn back on without problem
14
24/08/201210:00:0024/08/201210:00:002ONInstalled shileding on laser head
15
21/08/201214:00:0022/08/201223:00:003ONB field 0T
16
16/08/201217:00:0016/08/201218:00:004OFFPC08 rebooted. After that, USB error: no more shutter. Need to run special command to enable USB driver (first time): "chmod o+rw /dev/ttyUSB0"
17
11/08/2012-14/08/201208:00:003ONB field 0T
18
02/08/201214:00:0002/08/201216:00:005OFFPower measurement. filter damaged
19
11/07/201214:30:0011/07/201217:00:005OFFPower measurement
20
25/06/201209:00:0029/06/201212:00:006ON/OFFtechnical stop. Liyuan  . new mirror. +2 degrees room temp.
21
19/06/201206:00:0028/06/201217:00:003ON/OFFB field 0T
22
12/06/201217:00:0012/06/201217:30:006OFFhalf of the power : internal attenuation wrong position.
23
11/06/201209:00:0011/06/201212:00:001OFFreplacement of the chiller following Ta instabilities (temp water)
24
06/06/201213:02:0006/06/201214:30:004OFFShutter error
25
26
27
25/04/2012-27/04/2012-6OFFTS. power measurement : clean fibre -> change attenuation from 100 to 40% because of too high level
28
18/04/2012-18/04/2012-7ONchange current from 45 to 55 A.
29
30
31
32
33
34
35
36
37
38
39
40
41
42
43
44
45
46
47
48
49
50
51
52
53
54
55
56
57
58
59
60
61
62
63
64
65
66
67
68
69
70
71
72
73
74
75
76
77
78
79
80
81
82
83
84
85
86
87
88
89
90
91
92
93
94
95
96
97
98
99
100
Loading...
Main menu